OpenStudy (anonymous):

For the function of f and g, find the following and state the domain of each result f(x) = 5x+2/3x-4 ; g(x)= 2x/3x-4. How do I solve this pls...

jhonyy9 (jhonyy9):

first of all when is a fraction undefined ?

OpenStudy (anonymous):

when the denominator is equal to zero i think

jhonyy9 (jhonyy9):

yes right so than this will be the first you need calculi for what value will be the denominator zero ok ?

jhonyy9 (jhonyy9):

so what will be the restriction of x ?

OpenStudy (anonymous):

3x-4 = 0 3x = 4 x=4/3 is this right?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

what is the restriction on x? x is not equal to 0?

jhonyy9 (jhonyy9):

x not equal 4/3 never because than these fractions will be undefined yes ?

jhonyy9 (jhonyy9):

because x=4/3 result denominator equal zero

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so what is the domain for f(x)?

jhonyy9 (jhonyy9):

moment now you need make f(x) equal zero and g(x)=0 and calculi ok ?

OpenStudy (anonymous):

ok working on it....

jhonyy9 (jhonyy9):

try calculi you here ,ok ?

jhonyy9 (jhonyy9):

courage please i will help you

OpenStudy (anonymous):

so will it be 5(4/3) + 2 divided by 3(4/3) - 4 = 0? is this right?

OpenStudy (anonymous):

i got 7/-1

jhonyy9 (jhonyy9):

5x+2 ------ = 0 3x-4 5x+2=0 5x=-2 x= -2/5 how you have got this result ?

OpenStudy (anonymous):

I plugged in 4/3 for x. guess i was wrong.

jhonyy9 (jhonyy9):

x=4/3 is a restriction of x

OpenStudy (anonymous):

ok so we exclude that right?

jhonyy9 (jhonyy9):

yes sure

OpenStudy (anonymous):

ok am solving for g(x)

jhonyy9 (jhonyy9):

let s go but here ok ?

OpenStudy (anonymous):

2x=0 x=-2

jhonyy9 (jhonyy9):

2x=0 so try prove it for x=-2 you will get zero ? i think never

jhonyy9 (jhonyy9):

how you have got this -2 ?

jhonyy9 (jhonyy9):

so there is 2x=0 divide both sides by 2 what will get ?

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

sorry... 2x=0 x = 0

jhonyy9 (jhonyy9):

yes sure this is easy yes ?

OpenStudy (anonymous):

yayyy got it!... with your help yes it was easy.. now how to write the domain...

jhonyy9 (jhonyy9):

so how you know it where is defined a function f(x) or g(x) so when these are fractions ?

jhonyy9 (jhonyy9):

do you have learned it in your school,yes ?

OpenStudy (anonymous):

yes but online and it is so confusing to follow

jhonyy9 (jhonyy9):

are defined in R without this restriction what mean R ?

OpenStudy (anonymous):

what is R?

jhonyy9 (jhonyy9):

you dont know it ?

OpenStudy (anonymous):

no.

jhonyy9 (jhonyy9):

how you not the set of real numbers ?

OpenStudy (anonymous):

ok sure the symbol for the set of all real numbers! yes i do

OpenStudy (anonymous):

I usually write it as x|x

jhonyy9 (jhonyy9):

yes so than you know what is the answer on this exercise now ?

OpenStudy (anonymous):

I guess it is the x|x >=4/3 for f(x) and g(x) x|x>= 2

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

is that wrong? pls correct me

jhonyy9 (jhonyy9):

so the domain will be R without this restriction of x

OpenStudy (anonymous):

so will that be x|x not equal to 4/3?

OpenStudy (anonymous):

hiya Jhonyy9 are you there?

jhonyy9 (jhonyy9):

yes i m here

jhonyy9 (jhonyy9):

the domain is R without 4/3 ok ?

OpenStudy (anonymous):

I got it. thanks so much. You have been very helpful.

jhonyy9 (jhonyy9):

np was my pleasure good luck